OpenStudy (anonymous):

A numeric code consists of four digits. The first digit must be less than 5 and greater than 0 (zero), but the other digits can be any numeral from 0 to 9. How many different codes are possible?

OpenStudy (bahrom7893):

0,1,2,3,4 => 5 0, 1, 2, 3, 4, 5, 6, 7, 8, 9 => 10 5*10*10*10 = 5,000 codes

OpenStudy (anonymous):

10^3 codes are possible with the last 3 digits. As there are 4 numbers (1, 2, 3, 4) greater than 0 but smaller than 5, the total number of codes is 4*10^3=4*1000=4000

OpenStudy (bahrom7893):

ohh cheese is right

OpenStudy (bahrom7893):

Silly me, i didn't read and greater than 0 part.